# Interplay between strain and size quantization in a class of topological insulators based on inverted-band semiconductors

### v1

@@ VII Conclusions
We have considered several problems in typical experimental situations related to topological surface states when a sample of the inverted-band material has a finite size and is subjected to strain. The main interest is to understand the physics of the series of transitions which happens at a given strain value with changing a sample width. In particular, we have considered analytically the DSM -3D quantized topological insulator transition. It is believed to occur in the case when for a large sample width one has a compressive in-plane strain which leads to an overlap of the light and heavy-hole bands (DSM regime). For this case our conclusion is actually opposite to those reached by the majority of the researchers. We have shown that near the transition point the surface state (DK1) cannot be treated as a truly topological one. At least this state cannot exhibit the topological properties in any transport experiments. The parameters of the problem near this point are such that an appreciable overlap of the surface states’ wave functions located at opposite boundaries occur. As a result, an elastic (and spin conserving) impurity scattering between the states located at opposite boundaries will destroy the robustness of the surface state.
@@
While solving this problem we have obtained several other interesting and important results. For example, the effective boundary conditions for the solutions of the Kane model (with the contributions from the far bands) in the case of high barriers allows the use of this model to solve analytically various finite-size problems in the presence of strain for a realistic situation when heavy holes have finite masses. We hope that the results obtained can help in interpretation of current and future experimental data.
@@
Appendix A Derivation of Eq. (41)
@@
We derive here Eq. (41) and also obtain a shift of the central point of the V-P branches due to finite size effects in the limit of large width ($\alpha/s\gg 1$). We rewrite Eq. (39) in the following form
@@
The first and the last terms determine the gap value, and the second one gives the shift of the central point of the V-P branches
@@
Appendix B Exact equation for $k_{y}^{\star}$ obtained within the Luttinger model for a finite-size sample in the presence of strain
@@
Note that in this Appendix the definition of $\gamma,\tilde{\gamma}$ is different from the one in the main text. Here these constants are ”full” Luttinger parameters, thus $02\tilde{\gamma}-\gamma>0,2\tilde{\gamma}+\gamma>$ 0. Eq. (48) is obtained from the condition $\kappa_{l}=0$. The expression for $\kappa_{h}$ presented above follows from the equation

### v2

@@ VIII Conclusions
We have considered several problems in typical experimental situations related to topological surface states when a sample of the inverted-band material has a finite size and is subjected to strain. The main interest is to understand the physics of the series of transitions which happens at a given strain value with changing a sample width. In particular, we have considered analytically the DSM -3D quantized topological insulator transition. It is believed to occur in the case when for a large sample width one has a compressive in-plane strain which leads to an overlap of the light and heavy-hole bands (DSM regime). We have identified all the bulk-like and surface-like subbands which are involved in the transition, and, in particular, have shown that the branch of the surface states which is involved is the DK1 states (in strong contrast to the classification of the states used by the other researchers Ohtsubo ; Coster ; Anh ; Vail ). We have made concrete predictions about the robustness of the topological states near the transition point. This conclusion opposes those reached by the majority of the researchers. We have shown that near the transition point the surface state (DK1) cannot be treated as a truly topological one. At least this state cannot exhibit the topological properties in any transport experiments. The parameters of the problem near this point are such that an appreciable overlap of the surface states’ wave functions located at opposite boundaries occurs. As a result, an elastic (and spin-conserving) impurity scattering between the states located at opposite boundaries will destroy the robustness of the surface state.
We have also proposed a possible qualitative explanation of the discrepancy between the existing experimental ARPES data and most theoretical predictions regarding the topological surface states spectra which is based on the pattern of a bending of the energy bands of the topological material near the interface with the direct-band material.
@@
While solving these problems we have obtained several other interesting and important results. For example, we have derived the effective boundary conditions for the solutions of the Kane model (with the contributions from the far bands) in the case of high barriers. It allows the use of this model to solve analytically various finite-size problems in the presence of strain for a realistic situation when heavy holes have finite masses. We hope that the results obtained can help in interpretation of current and future experimental data.
@@
Appendix A Boundary conditions
A.1 High-energy barriers between inverted and direct materials
The wave functions (spinors) in all regions have the form
Here $u$ corresponds to the electron contribution, while $v$ and $w$ describe the light-hole and heavy-hole contributions, respectively. We consider everywhere (except section IV) the case of high-energy offsets in the conduction and valence bands when the energy gap of the direct barrier material is large compared to the energy gap of the inverted-band material. Using this condition we find the following decaying solutions in the barrier region, $x>0$, for the heavy and light particles:
$\Psi_{h}=e^{-\kappa_{hb}x}\left(\begin{array}[]{ll}0&\\ 0&\\ 1\end{array}\right),\Psi_{l}=e^{-\kappa_{lb}x}\left(\begin{array}[]{cc}-i\sqrt{\frac{1-\eta}{\eta}}&\\ 1&\\ 0\end{array}\right)$ (35)
Function Eq.(28) in the barrier region is a linear superposition of the functions Eq.(35). Here $\kappa_{lb}=\sqrt{3\eta(1-\eta)/2}E_{gb}/P$, and $\kappa_{hb}=\sqrt{(1-\eta)E_{gb}/(\gamma-2\tilde{\gamma})_{b}}$. The parameter $\eta$ ($10<\eta<$ 1) describes the asymmetry of the offsets in the conduction and valence bands, see Fig. 1. The edges of the conduction and valence bands in the barrier region are $E_{cb}=\eta E_{gb}$ and $E_{vb}=(\eta-1)E_{gb}$, where $E_{gb}>0$ is the energy gap of the direct-gap (barrier) material. The zero of the energy coincides with the middle of the $\Gamma_{8}-\Gamma_{6}$ gap of the inverted material. The case $\eta=1/2$ corresponds to symmetric offsets.
A.2 Effective boundary conditions
We start with the derivation of the boundary conditions (BCs) applicable for the case of arbitrary offsets in the conduction and valence bands. We use the standard procedure, i.e. we integrate all components of the Schrödinger equation with the Hamiltonian Eq. (2) across the boundary $x=0$ from region 1 to region 2. The terms containing the $\gamma$ and $\tilde{\gamma}$ we symmetrize in the standard manner, see Eq. (20) in Ch. III, Sec. II of Ref. Bastard, . For example, the term $-2i\sqrt{3}k_{y}\tilde{\gamma}\hat{k}_{x}w(x)$ is presented as $-i\sqrt{3}k_{y}(\tilde{\gamma}\hat{k}_{x}w(x)+\hat{k}_{x}(\tilde{\gamma}w(x)))$, etc. Then after intergation of the second and third equations we obtain
$\frac{-2iP}{\sqrt{6}}(u_{2}-u_{1})+\gamma_{lb}v_{2}^{\prime}-\gamma_{l}v_{1}^{\prime}$ $=$ $\sqrt{3}k_{y}\delta\tilde{\gamma}w_{1},$
$\sqrt{3}k_{y}\delta\tilde{\gamma}v_{1}+\gamma_{hb}w_{2}^{\prime}-\gamma_{h}w_{1}^{\prime}=0,$ (36)
where $\delta\tilde{\gamma}=\tilde{\gamma}_{b}-\tilde{\gamma}$, $\gamma_{hb}=(\gamma-2\tilde{\gamma})_{b}$, etc., and we have already used that $v_{1}=v_{2}$, $w_{1}=w_{2}$. The prime means the first derivative and indexes $1,2$ refer to the well region 1 and the barrier region 2. From the explicit form of the barrier functions Eq. (35), we have the following equalities at $x=0$: $u_{2}=-iv_{1}\sqrt{(1-\eta)/\eta}$, $v_{2}^{\prime}=-\kappa_{lb}v_{1}$, $w_{2}^{\prime}=-\kappa_{hb}w_{1}$, where again we used the continuity of $v,w$ components. Then we obtain the effective BCs which contain only the quantities of region 1
$\frac{-2iP}{\sqrt{6}}(-i\sqrt{\frac{1-\eta}{\eta}}v_{1}-u_{1})-\gamma_{lb}\kappa_{lb}v_{1}-\gamma_{l}v_{1}^{\prime}$ $=$ $\sqrt{3}k_{y}\delta\tilde{\gamma}w_{1},$
$\sqrt{3}k_{y}\delta\tilde{\gamma}v_{1}-\gamma_{hb}\kappa_{hb}w_{1}-\gamma_{h}w_{1}^{\prime}=0.$ (37)
We assume the condition $\gamma,\tilde{\gamma}\ll P^{2}/E_{g}$, where $E_{g}>0$ is the $\Gamma_{8}-\Gamma_{6}$ energy gap in the inverted-band material, and similar condition in the barrier region. The functions $v_{1},w_{1}$ in the region 1 have the form
$v_{1}(x)=Ae^{\kappa_{l}x}a_{2}+Be^{\kappa_{h}x}b_{2},$
$w_{1}(x)=Ae^{\kappa_{l}x}a_{3}+Be^{\kappa_{h}x}b_{3}$ (38)
Assuming in addition the high barrier conditions $k_{y},\kappa_{l},\kappa_{h}\ll\kappa_{lb},\kappa_{hb}$, one obtains the following hard-wall effective boundary conditions for the wave function inside the gapless material 1, see Fig. 1:
$\left(u+iv\sqrt{\frac{1-\eta}{\eta}}\right)|_{x=0}=0;\,\,\,w|_{x=0}=0$ (39)
We note that at the opposite boundary ($x=-d$) the sign in front of the imaginary unity in Eq. (39) should be changed for the negative one.
In real structures the values of $\gamma$ (or $\tilde{\gamma}$) parameters are of the same order of magnitude in both regions 1 and barrier region 2, i.e. $\gamma_{l}\simeq\gamma_{lb}$, for example. That is why the condition of the applicability of Eq. (39) is
$E_{g}\ll E_{gb}\ll P^{2}/\gamma$ (40)
By the effective boundary conditions we mean the following. We can consider the solution of the Schrodinger equation only in the inner region 1, all the information about the barrier regions enters the problem through the BCs presented by Eq. (39). Moreover, in the high barrier limit we avoid the usual troubles related to the choice of the order of different noncommuting operators in the Hamiltonian, Eq. (2). For example, in the popular book by G. Bastard Bastard the author uses the very well-known BCs (a wave function and the current flux continuities) for the case of the finite offsets values. In this case one needs to take care about the order of the non-commuting operators in the Hamiltonian. G. Bastard uses the anti-commutator of the operators to satisfy the hermiticity of the Hamiltonian. However, it is the well-known fact that this combination is not unique, see also Ref. Burt, . We should mention also that we are not aware of any work where the boundary conditions Eq. (39) were obtained.
It is remarkable that even in the limit of very high barriers the BCs do not reduce to the trivial zero form ($u=0,v=0,w=0$), but contain the information about the asymmetry of the offsets in the conduction and valence bands. It is also important to mention that while considering the surface states problem with high-energy barriers and for energies near the $\Gamma_{8}$ point (i.e. much smaller than the energy gap $E_{g}$) one can neglect the contribution from the electron $\Gamma_{6}$ band. Then the boundary conditions within the Luttinger model reduce to $v=0,w=0$, i.e. to zero values of the light-hole and heavy-hole components Dyak ; Dyak1 .
To assess how realistic the approximation of the high barrier is for the real material CdTe/HgTe, we can mention the following. The general condition of the high barrier requires the following inequalities, $\eta E_{gb}\gg E_{g},(1-\eta)E_{gb}\gg E_{g}$, see Fig. 1. It means that the decaying factor $\kappa$ of the wave function of the light particles under the barrier should be much larger than in the well region (HgTe). (For the heavy holes this factor is much larger than for the light particles). This quantity $\kappa$ in the material region has totally different value for the VP surface states and for the DK states.
In the case of the VP states we can estimate the location of the Dirac point of the VP states, Eq. (17), in two cases: for the infinite barrier and for the real parameters of the CdTe/HgTe structure and zero strain. For the real structure one has $\epsilon_{DP}$=61 meV compared to $\epsilon_{DP}$=72 meV in the infinite barrier case. We have used $\eta$ =0.74. Thus it makes an error of less than 4 % for the position of the Dirac point on the scale of $E_{g}$. In absolute units, the error amounts to an energy shift of about 10 meV, for surface states buried by $\sim$100 meV under the closest bulk band edge.
As for the DK states, which play the major role in the descriptions of physics in III and VI sections, the situation is much better. The reason is that the energy of the DK1 state in the HgTe region is close to the edge of the $\Gamma_{8}$- band, and the corresponding factor $\kappa$ for that region can be arbitrary small (depending on the $k_{y}$ value).
Appendix B Correction to Eq. (17) due to finite values of $\gamma$ and $\tilde{\gamma}$
Let us derive the correction to Eq. (17) due to interaction with the far bands, i.e. we will take into account the terms proportional to $\gamma$ and $\tilde{\gamma}$ in the Hamiltonian, Eq. (2). For $k_{y}=0$ the heavy holes are completely decoupled from the light particles, and to find the location of the DP of the Volkov-Pankratov states we need to solve the 2x2 problem (the top left corner of the matrix Eq. (2)). The wave function is a two-component spinor
$\psi=\left(\begin{array}[]{ll}u(x)&\\ v(x)\end{array}\right)$ (43)
To obtain the boundary conditions (BCs) we use the standard procedure, i.e. we integrate the both components of the Schrödinger equation across the boundary $x=0$. The term $\gamma_{l}\hat{k}_{x}^{2}$ ($\gamma_{l}=\gamma+2\tilde{\gamma}$) we symmetrize in the standard manner: Bastard $-\partial/\partial x\gamma_{l}\partial/\partial x$. Then we obtain the following BCs
$\frac{-2iP}{\sqrt{6}}(u_{2}-u_{1})+\gamma_{l}^{(2)}v_{2}^{\prime}-\gamma_{l}^{(1)}v_{1}^{\prime}=0,$
$v_{1}=v_{2},$ (44)
where the prime means the first derivative and indexes $1,2$ refer to the regions $1,2$. The wave functions decay in the left region 1 as $\exp(\kappa_{1}x)$ and in the right region 2 as $\exp(-\kappa_{2}x)$, with positive $\kappa_{1,2}$. The values of these quantities are (for $\Delta=0$)
$\kappa_{i}^{2}=\frac{(\epsilon_{ci}-\epsilon)(\epsilon-\epsilon_{vi})}{2P^{2}/3+\gamma_{l}^{(i)}(\epsilon_{ci}-\epsilon)};i=1,2$ (45)
Using the BCs Eq. (44) we finally obtain
$\frac{2}{3}P^{2}\left[\kappa_{1}+\kappa_{2}\frac{(\epsilon-\epsilon_{c1})}{(\epsilon-\epsilon_{c2})}\right]=(\epsilon-\epsilon_{c1})(\kappa_{1}\gamma_{l}^{(1)}+\kappa_{2}\gamma_{l}^{(2)})$ (46)
Eq. (45) and Eq. (46) should be solved together to find the wanted energy $\epsilon$. In the first order with respect to small quantities $\gamma_{l}^{(1,2)}$, writing $\epsilon=\epsilon_{0}+\delta\epsilon$, one obtains
$\delta\epsilon=\frac{1}{\left[1-\frac{\epsilon_{0}(\epsilon_{0}-\epsilon_{c2})}{\kappa_{10}\kappa_{20}2P^{2}/3}\right]}\frac{(\kappa_{10}\gamma_{l}^{(1)}+\kappa_{20}\gamma_{l}^{(2)})}{(\kappa_{10}+\kappa_{20})4P^{2}/3}(\epsilon_{0}-\epsilon_{c1})(\epsilon_{0}-\epsilon_{c2})$ (47)
Here
$\kappa_{10}^{2}=\frac{(\epsilon_{c1}-\epsilon_{0})(\epsilon_{0}-\epsilon_{v1})}{2P^{2}/3},\,\kappa_{20}^{2}=\frac{(\epsilon_{c2}-\epsilon_{0})(\epsilon_{0}-\epsilon_{v2})}{2P^{2}/3},$
and $\epsilon_{0}$ is given by Eq. (17) for $\Delta=0$. Shift $\delta\epsilon$ is negative for both cases we consider in this work (HgTe/CdTe and $\alpha$-Sn/InSb). To illustrate Eq. (47) we present here the answer for the case of the structure with the symmetric offsets ($\eta=1/2$) when $\epsilon_{0}=0$
$\delta\epsilon=-\frac{3(E_{g}\gamma_{l}^{(1)}+E_{gb}\gamma_{l}^{(2)})}{16P^{2}}\frac{E_{g}E_{gb}}{E_{g}+E_{gb}}$ (48)
Note that here the following parameters are assumed to be small: $\gamma_{l}^{(1)}E_{g}/P^{2}\ll 1$, $\gamma_{l}^{(2)}E_{gb}/P^{2}\ll 1$.
Appendix C Derivation of Eq. (21)
@@ A.2 Effective boundary conditions
We derive here Eq. (21) and also obtain a shift of the central point of the VP branches due to finite size effects in the limit of large width ($\alpha/s\gg 1$). We rewrite Eq. (19) in the following form
@@
The first and the last terms determine the gap value, and the second one gives the shift of the central point of the VP branches
@@
Appendix D Exact equation for $k_{y}^{\star}$ obtained within the Luttinger model for a finite-size sample in the presence of strain
@@
Note that in this Appendix the definition of $\gamma,\tilde{\gamma}$ is different from the one in the rest of the paper. Here these constants are ”full” Luttinger parameters, thus $02\tilde{\gamma}-\gamma>0,2\tilde{\gamma}+\gamma>$ 0. Eq. (52) is obtained from the condition $\kappa_{l}=0$. The expression for $\kappa_{h}$ presented above follows from the equation